Can someone help me on this problem?
I need to solve: -6(x + 3) -10 = 3(-2x -5) -13
Thanks, Lisa

Answer by funmath(2873)

Hi Lisa,
First simplify both sides of the equal sign, then try to get all the x's on one side and those without x's on the other within the rules of algebra. Whatever you do to one side you must do to the other sides.
-6(x + 3) -10 = 3(-2x -5) -13 Simplify by Multilying first.
-6x-18-10=-6x-15-13 Simplify by combining like terms
-6x-28=-6x-28
Hey! Both sides are the same, so this is an identity. The solution is all real numbers. No matter what number you put in for x, both sides will be equal. If you keep trying to solve it you'll get 0=0:
6x-6x-28=6x-6x-28
0-28=0-28
-28=-28
28-28=28-28
0=0
Happy Calculating, Lisa!!!